jigglypuff314 (jigglypuff314):

If y = 2x - 8 , what is the minimum value of the product xy? (A) -16 (B) -8 (C) -4 (D) 0 (E) 2

jigglypuff314 (jigglypuff314):

If you're not too busy, could you help me please? @TuringTest @sourwing

OpenStudy (anonymous):

minize x(2x-8)

jigglypuff314 (jigglypuff314):

what's that mean?

OpenStudy (anonymous):

i meant minimize x(2x-8). Should be at the vertex

jigglypuff314 (jigglypuff314):

how do I minimize? is it like find the derivative?

OpenStudy (anonymous):

yes. But since this is a quadratic that opens up, you know the minimum is at the vertex

jigglypuff314 (jigglypuff314):

mmm but how did you get x(2x-8) if the given equation was y = 2x - 8 ?

jigglypuff314 (jigglypuff314):

oh wait, I see it nvm

OpenStudy (anonymous):

ok

jigglypuff314 (jigglypuff314):

so d/dx (x(2x-8)) = 4x - 8 and 4x-8=0 --> x = 2 ?

OpenStudy (anonymous):

yes, what's y?

OpenStudy (anonymous):

and what is xy?

jigglypuff314 (jigglypuff314):

mmm y = 2(2) - 8 = -4 so xy = (2)(-4) = -8 ok got it now :) Thank You!